# AutoGuide: Automated Generation and Selection of State-Aware Guidelines for Large Language Model Agents

### v1

@@ 5 Conclusion
In this paper, we presented AutoGuide, an effective framework for extracting and exploiting important domain knowledge from offline experiences for improving decision making with pre-trained LLMs. Pre-trained LLMs often lack the knowledge required for solving tasks in downstream domains, and to fill in such knowledge gap, we proposed to generate state-aware guidelines that can be incorporated into prompts for LLM agents. As AutoGuide extracts the guidelines by contrasting successful and failed trajectories on the same task, the resulting state-aware guidelines carry critical information for preventing failures in the domains. For inference, it provides the guidelines pertinent to each of the different states that LLM agents encounter, which can make pre-trained LLMs strong decision-making agents in the downstream domains. Empirically, we showed that AutoGuide outperforms strong LLM-based baselines by a large margin and achieves outstanding performance in multiple sequential decision-making benchmarks.
Appendix A Evaluation Details
A.1 ALFWorld (Shridhar et al., 2021)
A.1.1 Environment details
@@
Alternatively, the agent can generate think actions for planning and reflection, which helps with decision-making but does not change the environment itself. After one action is performed, the environment returns an observation that describes view changes.
@@
A.1.2 Baseline and Models
For ALFWorld tasks, we follow the same setting as ReAct by providing 2 in-context examples for each of the 6 task categories. The original results of ReAct in their paper are produced based on Text-Davinci-002. However, this GPT version is no longer available, so we apply gpt-3.5-turbo-instruct instead to generate actions. For ExpeL, we directly take the guidelines from their appendix and append them to the ReAct agent at test time.
A.1.3 Implementation details of AutoGuide
We run the first 100 training tasks of ALFWorld to collect $(\bm{\tau}_{\text{success}},\bm{\tau}_{\text{fail}})$ pairs with ReAct+Reflexion and extract state-dependant guidelines on the collected data. For state summarization, we provide 2-shot demonstrations for each of the 6 task categories. The corresponding prompt templates can be found in appendix B. All parameter details are shown in Table 5.
@@ A.1.3 Implementation details of AutoGuide
State Summarization Model gpt-3.5-turbo-instruct
@@
A.2 WebShop (Yao et al., 2022a)
A.2.1 Environment details
@@
A.2.2 Baseline and Models
Following ReAct, experiments are done in a one-shot setting. We apply gpt-3.5-turbo-0613 to generate actions, but when the token number exceeds the token limit (for example, for the n-shot ReAct experiments in Table 3), we use the 16k version of gpt-3.5-turbo-0613 instead. For ExpeL, we could not find how many training tasks the framework used for training. Therefore, we directly apply the guidelines from the appendix of their paper at test time. We only consider ExpeL with guidelines, not ExpeL with in-context example selection in our experiments for a fair comparison. The in-context example selection method is orthogonal to our work and can be easily combined with our method. For Reflexion, as shown in their paper, their 2-shot Reflexion prompt does not work well on WebShop. Therefore, we re-write a zero-shot prompt and apply gpt-4-1106-preview to generate episode-level reflections for all Reflexion experiments. Following Reflexion and ExpeL, the evaluation is done on 100 test tasks. The maximum number of allowed actions for each task is 15. At the same time, each search action shows the top 3 products for the search query. Please refer to Table 6 for more details about the experiments.
A.2.3 Implementation details of AutoGuide
We randomly sample 50 training tasks from the training set of WebShop, on which we run ReAct+Reflexion to collect pairs and generate guidelines. The state summarization prompt is one-shot, which is shown in appendix B. At test time, we ask gpt-3.5-turbo-0613 to select the most relevant top 2 guidelines for each state.
@@ A.2.3 Implementation details of AutoGuide
State Summarization Model gpt-3.5-turbo-0613
@@
A.3 WebArena (Zhou et al., 2023)
A.3.1 Environment details
@@
  • type[text][1 for enter or 0 for not enter]
@@ A.3.2 Baseline and Models
A.3.3 Implementation details of AutoGuide
Among the 19 human demonstrations provided in WebArena, 17 of them successfully complete the tasks. We directly run ReAct on the successful tasks to collect failure actions and generate guidelines correspondingly. We take advantage of the URL and webpage titles, which are a part of the observation for this benchmark, to help with state summarization.
@@ A.3.3 Implementation details of AutoGuide
State Summarization Model gpt-4-1106-preview
@@
Appendix B Prompt Templates
B.1 State Summarization
@@
Figure 6: Our prompt template for state summarization (Equation 2 from Section 3.3) in the WebShop domain.
@@
Figure 7: Our prompt template for state summarization (Equation 2 from Section 3.3) in the ALFWorld domain.
@@
Figure 8: Our prompt template for state summarization (Equation 2 from Section 3.3) in the WebArena domain.
@@
We present our prompt templates for state summarization with $\mathcal{M}_{\text{state}}$ (Equation 2 from Section 3.3) for WebShop, ALFWorld, and WebArena in Figures 6, LABEL:, 7, LABEL: and 8, respectively. In ALFWorld, there exist six categories of tasks, and we use state summarization prompting with 2-shot examples for each task, following the practice by Yao et al. (2022b). Figure 7 shows one example for the pick_and_place tasks.
B.2 Guideline Extraction
@@
In Figures 9, LABEL:, 10, LABEL: and 11, we state the prompt templates we use with $\mathcal{M}_{\text{guideline}}$ for extracting guidelines (Equation 3 from Section 3.3) in the WebShop, ALFWorld, and WebArena domains.
@@
Figure 9: Our prompt template for guideline extraction (Equation 3 from Section 3.3) in the WebShop domain.
@@
Figure 10: Our prompt template for guideline extraction (Equation 3 from Section 3.3) in the ALFWorld domain.
@@
Figure 11: Our prompt template for guideline extraction (Equation 3 from Section 3.3) in the WebArena domain.
B.3 State Matching
@@
Figure 12: Our prompt template for state matching (Sections 3.3, LABEL: and 3.4) in all the WebShop, ALFWorld, and WebArena domains.
@@
Figure 12 shows our prompt template for matching the summarized state with one of the existing states if there is any similar state, for the construction of the set of state-aware guidelines (Section 3.3) and the retrieval of relevant guidelines during testing (Section 3.4) in all the three domains: WebShop, ALFWorld, and WebArena.
B.4 Guideline Selection
@@
Figure 13: Our prompt template for selecting most relevant state-aware guidelines during the test time (Equation 4 from Section 3.4) in the WebShop and ALFWorld domains.
@@
Figure 14: Our prompt template for selecting most relevant state-aware guidelines during the test time (Equation 4 from Section 3.4) in the WebArena domain.
@@
For selecting only $k$ most relevant guidelines in case there are more corresponding state-aware guidelines during testing (Equation 4 from Section 3.4), we use the prompt with Figure 13 for WebShop and ALFWorld and Figure 14 for WebArena.
Appendix C Example State-Aware Guidelines
In Figure 15, we show a list of possible states and state-aware guidelines on WebArena.
@@ Appendix C Example State-Aware Guidelines
Figure 15: Example states and corresponding guidelines for WebArena.

### v2

@@ 5 Conclusion
We present AutoGuide, an effective framework for exploiting important domain knowledge from offline experiences for improving decision-making with pre-trained LLMs. We proposed to generate context-aware guidelines that can be incorporated into prompts for LLM agents. As AutoGuide extracts the guidelines by contrasting trajectories in offline data, the resulting context-aware guidelines carry critical information for preventing failures in the domains. For inference, it provides the guidelines pertinent to each of the different context that LLM agents encounter, which can make pre-trained LLMs strong decision-making agents in the downstream domains. Empirically, we showed that AutoGuide outperforms strong baselines by a large margin and achieves outstanding performance in decision-making benchmarks.
6 Acknowledgements
This work was supported in part by LG AI Research.
Appendix A Limitation and Broader Impacts
Limitation. The performance of AutoGuide depends on the diversity of offline experiences. As such, one important direction for improvement is to automatically collect diverse offline experiences through continual learning, where we iteratively generate guidelines and use them to gather more trajectories with high rewards. Another avenue is the need for quantifying the quality of generated contexts and guidelines. Currently, apart from applying context-aware guidelines to ReAct and measuring test time performance, there lacks a standardized method for quantifying the quality of generated contexts and selected guidelines. Introducing a quantifiable metric to approximate the quality could pave the way for new optimization approaches such as reinforcement learning.
Broader impact. This paper introduces research aimed at enhancing the decision-making capabilities of LLMs. In terms of societal impact, while we develop a generic LLM-based autonomous agent, having biased offline datasets may lead to making decisions with suboptimal outcomes. Additionally, autonomous agents may be misused for malicious applications. To mitigate these risks, potential solutions would include diversifying datasets, implementing ethical oversight, ensuring transparency and accountability, engaging with stakeholders for a broader perspective, and incorporating security measures to prevent misuse. We believe that the research community, including ourselves, should responsibly advance LLM-based agent research, prioritizing societal well-being and ethical considerations.
Appendix B Evaluation Details
B.1 ALFWorld [24]
B.1.1 Environment details
@@
Alternatively, the agent can generate a think action for planning and reflection, which helps with decision-making but does not change the environment itself. After one action is performed, the environment returns an observation that describes view changes.
@@
B.1.2 Baseline and Models
For ALFWorld tasks, we follow the same setting as ReAct by providing 2 in-context examples for each of the 6 task categories. The original results of ReAct in their paper are produced based on Text-Davinci-002. However, this GPT version is no longer available, so we apply gpt-3.5-turbo-instruct instead to generate actions. For ExpeL, we directly take the guidelines from their appendix and append them to the ReAct agent at test time. For Reflexion, we generate reflective feedback based on the test rewards if the tasks fail. This reflective feedback is appended to the action generation prompt, immediately following the context-aware guidelines for AutoGuide + Reflexion. When there are multiple past episodes, we concatenate the reflection feedback from each episode. This process continues until the maximum number of episodes (or trials) is reached.
B.1.3 Implementation details of AutoGuide
We run the first 100 training tasks of ALFWorld to collect $(\bm{\tau}_{+},\bm{\tau}_{-})$ pairs with ReAct+Reflexion and extract context-dependant guidelines on the collected data. For context identification, we provide 2-shot demonstrations for each of the 6 task categories. The corresponding prompt templates can be found in appendix C. All parameter details are shown in table 7.
@@ B.1.3 Implementation details of AutoGuide
Context Identification Model gpt-3.5-turbo-instruct
@@
B.2 WebShop [33]
B.2.1 Environment details
@@
B.2.2 Baseline and Models
Following ReAct, experiments are done in a one-shot setting. We apply gpt-3.5-turbo-0613 to generate actions, but when the token number exceeds the token limit (for example, for the n-shot ReAct experiments in table 1), we use the 16k version of gpt-3.5-turbo-0613 instead. For ExpeL, we could not find how many training tasks the framework used for training. Therefore, we directly apply the guidelines from the appendix of their paper at test time. We only consider ExpeL with guidelines, not ExpeL with in-context example selection in our experiments for a fair comparison. The in-context example selection method is orthogonal to our work and can be easily combined with our method. For Reflexion, as shown in their paper, their 2-shot Reflexion prompt does not work well on WebShop. Therefore, we re-write the prompt and apply gpt-4-1106-preview to generate episode-level reflections for all Reflexion experiments. Same as the experiments in ALFWorld, the generated reflections are appended to the action generation prompt, immediately following the context-aware guidelines for AutoGuide + Reflexion. In cases where multiple past episodes are available, the reflection feedback from each episode is concatenated. This process is repeated until the maximum number of trials is reached. Following Reflexion and ExpeL, the evaluation is done on 100 test tasks. The maximum number of allowed actions for each task is 15. At the same time, each search action shows the top 3 products for the search query. Please refer to Table 8 for more details about the experiments.
B.2.3 Implementation details of AutoGuide
We randomly sample 50 training tasks from the training set of WebShop, on which we run ReAct+Reflexion to collect pairs and generate guidelines. The context identification prompt is one-shot, which is shown in Appendix C. At test time, we ask gpt-3.5-turbo-0613 to select each state’s most relevant top 2 guidelines.
@@ B.2.3 Implementation details of AutoGuide
Context Identification Model gpt-3.5-turbo-0613
@@
B.3 WebArena [8]
B.3.1 Environment details
@@
  • type [element_id] [text] [1 for enter or 0 for not enter]
@@ B.3.2 Baseline and Models
B.3.3 Implementation details of AutoGuide
We directly run ReAct on the tasks with $\bm{\tau}_{+}$ to collect $\bm{\tau}_{-}$ actions and generate guidelines correspondingly. We provide a 5-shot prompt for the context identification module, which is shown in Figure 7. At test time, the top 2 guidelines at each timestep are selected to guide action generation.
@@ B.3.3 Implementation details of AutoGuide
Context Identification Model gpt-4-1106-preview
@@
B.4 Real Websites
B.4.1 Environment details
We design a set of real-world website navigation tasks from 3 domains: Software Development (GitHub), Travel (Google Flights), and Education (Coursera), which have 30, 20, and 20 test tasks accordingly. Here are some example tasks:
  • GitHub
    • Navigate to the repository for the Python library Seaborn with the most stars and show me all the open issues labeled with bug.
    • Go to the GitHub org for Spotify and open the pinned project with the most stars for me.
  • Google Flights
    • Find me the one-way flight from Hong Kong to Paris departing on Oct 15th 2024 with the least emmisions.
    • Show me the booking options of the one-way flight departing from Auckland on September 11, 2024, and arriving in Rome with the earliest departure time on that day.
  • Coursera
    • Show me a Cybersecurity course that can finish within 1 month and show me all the reviews for the selected course.
    • Find me a Coursera guided project that covers Unity and show me its main page.
We follow the action space design of Visual WebArena [5], which has the following action types available:
  • goto [url]
  • click [element_id]
  • hover [element_id]
  • type [element_id] [text] [1 for enter or 0 for not enter]
  • press [key_combination]
  • scroll [up or down]
  • tab_focus [tab_index]
  • close_tab
  • go_back
  • go_forward
As the real websites are constantly and dynamically changing, we evaluate the completed task with human experts.
B.4.2 Baseline and Models
We directly run the two-shot SoM algorithm in the official codebase of Visual WebArena. The only modification we made from the original codebase is the bounding box detection algorithm, in which we further filter invisible bounding boxes from the list and add the list elements as interactable elements to consider.
B.4.3 Implementation details of AutoGuide
We provide a total of 6 training tasks (3 for GitHub, 2 for Google Travel, and 1 for Coursera), on which we collect human demonstration as $\bm{\tau}_{+}$’s and run SoM to collect $\bm{\tau}_{-}$’s. From the pairs with multi-modal observations we generate text-based guidelines to guide action selection. Both context identification and guideline extraction are done by gpt-4-vision-preview, and we provide a 3-shot prompt for context identification. All parameter details are shown in Table 10.
Parameter name Value
Allowed Episode Length 15
n-shots 2
Agent Model gpt-4-vision-preview
Context Identification Model gpt-4-vision-preview
Guideline Selection Model gpt-4-vision-preview
Guideline Extraction Model gpt-4-vision-preview
top-k guideline selection 2
Table 10: Experiment hyperparameters on multi-modal real-world website tasks.
Appendix C Prompt Templates
C.1 Context Identification
@@
Figure 5: Our prompt template for context identification (Equation 1) in the WebShop domain.
@@
Figure 6: Our prompt template for context identification (Equation 1) in the ALFWorld domain.
@@
Figure 7: Our prompt template for context identification (Equation 1) in the WebArena domain.
@@
We present our prompt templates for the context identification module $\mathcal{M}_{\mathrm{context}}$ (Equation 1) for WebShop, ALFWorld, and WebArena in Figures 5, 6 and 7, respectively. In ALFWorld, there exist six categories of tasks, and we use context identification prompting with 2-shot examples for each task, following the practice by [23]. Figure 6 shows one example for the pick_and_place tasks.
C.2 Guideline Extraction
@@
Figures 8, 9 and 10 detail our prompt templates $\mathcal{M}_{\mathrm{guideline}}$ for extracting guidelines (Equation 2) in the WebShop, ALFWorld, and WebArena domains.
@@
Figure 8: Our prompt template for guideline extraction (Equation 2) in the WebShop domain.
@@
Figure 9: Our prompt template for guideline extraction (Equation 2) in the ALFWorld domain.
@@
Figure 10: Our prompt template for guideline extraction (Equation 2) in the WebArena domain.
C.3 Context Matching
@@
Figure 11: Our prompt template for context matching (Sections 3.2 and 3.3) in all the WebShop, ALFWorld, and WebArena domains.
@@
Figure 11 shows our prompt template for matching the generated context with one of the existing contexts if there is any similar context, for the construction of the set of context-aware guidelines (Section 3.2) and the retrieval of relevant guidelines during testing (Section 3.3) in all the three domains: WebShop, ALFWorld, and WebArena.
C.4 Guideline Selection
@@
Figure 12: Our prompt template for selecting the most relevant context-aware guidelines during the test time (Equation 3 from Section 3.3) in the WebShop and ALFWorld domains.
@@
Figure 13: Our prompt template for selecting the most relevant context-aware guidelines during the test time (Equation 3) in the WebArena domain.
@@
For selecting only $k$ most relevant guidelines in case there are more corresponding context-aware guidelines during testing (Equation 3), we use the prompt with Figure 12 for WebShop and ALFWorld and Figure 13 for WebArena.
Appendix D Example Context-Aware Guidelines
In Figures 14 and 15, we show a list of possible contexts and context-aware guidelines on WebArena and real-world websites, respectively.
Figure 14: Example contexts and corresponding guidelines for WebArena.
@@ Appendix D Example Context-Aware Guidelines
Figure 15: Example contexts and corresponding guidelines for real-world websites.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: In the abstract and introduction, we clearly outline our main contribution: extracting context-aware guidelines from offline data to enhance the decision-making of LLM-based agents. We support the effectiveness of our method through empirical evaluations in Section 4.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We outline potential limitations of AutoGuide in Appendix A.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate ”Limitations” section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This paper does not include theoretical contributions.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Section 4.1 and Appendix B provide detailed information (e.g., model versions, hyperparameters, number of training data) for reproducing our experimental results. Additionally, Appendix C contains our prompt templates.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: We intend to release the source code upon acceptance.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We detail our experimental settings in Section 4.1 and Appendix B.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We provide the average rewards and success rates calculated across test tasks in our results, but we do not include the variance.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer ”Yes” if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Our experiments are based on GPT, and we detail the versions of GPT in Appendix B.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our work conforms the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We provide the broader impacts of our work in Appendix A.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper makes the algorithmic contribution and does not release new data or models.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We reference the original papers and sources throughout the paper.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: This paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.